A Hemispherical Dome,open at base,is made from a sheet of fibre.
if the radius of hemispherical dome is 40cm and 13/170 of the fibre sheet actually used was wasted in making the dome.
Find the cost of the dome at the rate of Rs.35 per cm square
(take π =3.14)

Answer by mananth(16946)

A Hemispherical Dome,open at base,is made from a sheet of fibre.
if the radius of hemispherical dome is 40cm and 13/170 of the fibre sheet actually used was wasted in making the dome.
Find the cost of the dome at the rate of Rs.35 per cm square
(take π =3.14)
Surface area of hemisphere = 2pi*r^2 r = radius
2*3.14*40*40 = 10048 cm square
Let the area of fibre sheet be x
10058+13/170 *x = x
1708160 +13x= 170x
1708160= 157x
= 10880 sq. cm
Rs.35/sq. cm = cost of fibre sheet.
10880 *35= Rs.380,800